OpenStudy (anonymous):

x - y + z = 5 3x + 2y - z = -2 2x + y + 3z = 10 (elimination or substitution is acceptable)

OpenStudy (anonymous):

1.add the first two eq. 2.multiply second by 3 then add in third

OpenStudy (texaschic101):

x - y + z = 5 3x + 2y - z = -2 --------------add 4x + y = 3 3x + 2y - z = -2 -->(3) 2x + y + 3z = 10 --------------- 9x + 6y - 3z = -6 (result of multiplying by 3) 2x + y + 3z = 10 ----------------add 11x + 7y = 4 4x + y = 3 --->(-7) 11x + 7y = 4 ----------------- -28x - 7y = -21(result of multiplying by -7) 11x + 7y = 4 -----------------add -17x = - 17 x = 1 4x + y = 3 4(1) + y = 3 4 + y = 3 y = 3 - 4 y = -1 x - y + z = 5 1 - (-1) + z = 5 1 + 1 + z = 5 z = 5 - 2 z = 3 check... 3x + 2y - z = -2 3(1) + 2(-1) - 3 = -2 3 - 2 - 3 = -2 -2 = -2 (correct) x = 1, y = -1, z = 3

OpenStudy (anonymous):

Didn't they already help you?

OpenStudy (anonymous):

yes but i noticed a part later that i didnt understand

OpenStudy (anonymous):

i dont understand where that -7 came from...

OpenStudy (anonymous):

I'm not good with this level of math, but maybe someone else can help you.

OpenStudy (anonymous):

okay, thank you anyway :)

OpenStudy (anonymous):

i dont see where that -7 came from...

OpenStudy (anonymous):

the idea is to make the coefficient of y or x the same. if they are of opposite sign ,then we add. if they are of same sign,then we subtract.